An office worker attempts to move a filing cabinet (m = 48 kg). The coefficient of static friction between the floor and the filing cabinet is 0.53.

a) Calculate the normal force acting on the filing cabinet.

b) Calculate the minimum horizontal force that the office worker must apply to the filing

cabinet to JUST move it.


1
Expert's answer
2021-01-21T09:56:09-0500

Friction force ="\mu*m*g"

Therefore frictional force=0.53*48*9.8 N = 249.312 N

Normal force = 48*9.8 = 470.4 N

To move filing, required force = frictional force to overcome = 249.312 N